OpenStudy (anonymous):

can someone help me find the inverse of f(x). and determine whether it is a function, and please state its domain and range.

OpenStudy (anonymous):

To find the inverse of a function, switch the y and x values, and then evaluate for y.

OpenStudy (anonymous):

\[f(x)=\sqrt[3]{x+1}\]

OpenStudy (anonymous):

So we'll start by swapping out y [f(x)] and x, then solving for y: \[x=(y+1)^{3}\] So to find y we'll cube root both sides: \[x^{1/3} = y+1\]And finally we'll subtract 1 from both sides: \[y=x^{1/3} -1\]

OpenStudy (anonymous):

so its a function with all real numbers ?

OpenStudy (anonymous):

Yes, that seems right to me, as you can take the cube root of any number x.

OpenStudy (anonymous):

alright thanks!(: